OpenStudy (anonymous):

2 QUICK QUESTIONS AND NEED HELP A.S.A.P 1.In the right triangle, a = 3 and b = 3. a.Find the exact value of c. Write the answer in simplest radical form. b.Approximate the value of c to the nearest hundredth.

OpenStudy (anonymous):

Kathy uses the Babylonian method to estimate 18 to the nearest hundredth. She starts with a guess of 4.5. a. Fill in the blanks for the first approximation. \[\frac{ ?+\frac{ ? }{ ? } }{ 2 }=?\] b. Fill in the blanks for the second approximation. Do not round any numbers until the end. \[\frac{ ?+\frac{ ? }{ ? } }{ 2 }\approx?\] c. Fill in the blanks for the third approximation. Do not round any numbers until the end. \[\frac{ ?+\frac{ ? }{ ? } }{ 2 }\approx?\]

OpenStudy (anonymous):

@ajprincess @jojo4eva @Loser66 @rose21 @Smores

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

sup

OpenStudy (anonymous):

can any on of yall hel me out

OpenStudy (anonymous):

help i meant

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my friend can help we just got in to this

OpenStudy (anonymous):

do you know anything about the problem how to start off

OpenStudy (phi):

I think they want you to use the pythagorean theorem

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@beastonastick and @phi can you guys come back and help me plz

OpenStudy (anonymous):

wht u need

OpenStudy (anonymous):

idk how to do this

OpenStudy (anonymous):

do wht

OpenStudy (anonymous):

the question that i posted

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@phi can you help me solve this but i mean step by step

OpenStudy (anonymous):

phi

OpenStudy (anonymous):

i have already seen that video

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

oh ok let me see it and see wht i can do wht grade are you in

OpenStudy (anonymous):

@phi can you come beck plz

OpenStudy (anonymous):

im in 8th grade

OpenStudy (anonymous):

i got somebody else who can help u

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@willis97

OpenStudy (anonymous):

hey @Nerdygeek1200 and @NaomiBell1997

OpenStudy (anonymous):

here he comes

OpenStudy (anonymous):

can you guys help me out here

OpenStudy (anonymous):

hey. Sorry i'm stuck on your problem, I can't figure it out, sorry :/ I tried, I really did

OpenStudy (anonymous):

exact answer is 3 radical 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

do you have a calculator

OpenStudy (anonymous):

i've got another question for you

OpenStudy (anonymous):

b. Approximate the value of c to the nearest hundredth

OpenStudy (anonymous):

4.24

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

can oyu give me a medal

OpenStudy (anonymous):

all of you guys plz